# Duality defect in a deformed transverse-field Ising model

### v3

@@ I Introduction
Compared with open boundaries, long-lived modes associated with defects are much less studied. Recently a special class of defects, the so-called topological defects, have attracted attention from the fields of condensed matter physics, high energy theory, and quantum information [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. In particular, Ref. [41] considered a 1d transverse field Ising (TFI) model on a circular chain with the insertion of a Kramers-Wannier [67] duality interface and a domain wall in the coupling space, bi-partitioning the chain into a ferromagnetic phase and a paramagnetic phase. In each $\mathbb{Z}_{2}$ charge ±$\pm 1$ sector of the system, there exists a zero mode supported on neighboring sites around the domain wall, which exactly commutes with the Hamiltonian. More recently, Ref. [55] studied the corresponding problem in a periodically-driven Floquet setting, and found an isolated zero mode localized around the domain wall, exactly commuting with the Floquet unitary. In the special limit where the duality interface is right next to the domain wall thereby shrinking one of the gapped phases to zero size, the zero mode becomes localized around the duality defect. Stability of the zero mode against deformations was probed in [68] for a particular deformed Floquet TFI model. In the presence of the deformation, the zero mode starts to decay. However, the decay mechanism is similar to a reservoir effect, such that for small system sizes the zero mode only partially decays. This can be quantified by the infinite-temperature auto-correlation function of a local operator overlapping with the zero mode, which stabilizes around a finite plateau at infinite time.
@@
Even though our methods apply to more general quantum systems, here we will streamline the ideas using the example of a 1d deformed TFI model in presence of a Kramers-Wannier (KW) duality defect, both in the Hamiltonian setting and in the periodically-driven Floquet setting. First, we derive the deformed defect Hamiltonian given in Eq. (6) for the general case using three different methods: half-chain KW transformation, utilization of techniques from the Ising fusion category, and defect-modified weak integrability breaking deformation. The results agree between these three quite distinct methods. We then work out the deformation of defect-modified higher charges in the model, and study their slower decay behavior under the time evolution by the deformed Hamiltonian. Finally, we move on to the corresponding deformed Floquet TFI model with a duality defect. The zero mode localized around the defect starts to decay in the presence of the deformation. Similar to [68], for small system sizes the zero mode only decays partially, as corroborated by a finite plateau of the infinite-temperature auto-correlation function for an overlapping local operator.
@@ II Transverse field Ising model with Kramers-Wannier self-dual deformation
The 1d TFI model can be deformed in different ways that are self-dual under the Kramers-Wannier transformation [69, 70]. Here we consider the model first studied in [69]. Our setup is a chain of $L$ sites with site-labels $j=1,...,L$. The Hamiltonian for the deformed TFI model with periodic boundary condition is given by
@@ III Duality defect in the self-dual deformed TFI model
Due to the KW self-dual nature of the deformation $H_{\lambda}$, we can construct a duality defect in this deformed TFI model. In later Sections, we will describe three different ways to do so. In Section IV, we construct the duality defect via half-space Kramers-Wannier transformation. In Section V, we illustrate the duality defect construction using techniques from fusion categories. In Section VI, we provide another way of deriving the duality defect Hamiltonian, as a special kind of weak integrability breaking deformation where the conserved charges in the undeformed model are modified by the defect.
@@ III.1 Fusion of two duality defects
In particular, the site $i_{0}-2$ only enters the transformed Hamiltonian through the last three terms in Eq. (12). The Pauli operator $Z_{i_{0}-2}$ commutes with the transformed Hamiltonian, consequently the whole defect Hilbert space admits a decomposition into the ±$\pm 1$ eigenspace of $Z_{i_{0}-2}$. In the $+1$ eigenspace, Eq. (12) reduces to the no-defect deformed TFI model, albeit on $L-1$ sites labeled by $1,\dots,i_{0}-3,i_{0}-1,\dots,L$. In the $-1$ eigenspace, we obtain the $\eta$-defect Hamiltonian (see Eq. (10)) in the deformed TFI model on $L-1$ sites. This defect Hilbert space decomposition, schematically expressed as
@@ IV Duality defect from half-chain Kramers-Wannier transformation
$\begin{split}\widehat{D}X_{2r+1}&=\widetilde{Z}_{2r}\widetilde{Z}_{2r+2}\widehat{D}~{},\\ \widehat{D}Z_{2r-1}Z_{2r+1}&=\widetilde{X}_{2r}\widehat{D}~{}.\end{split}$ (16)
@@
The action of $\widehat{D}_{2i_{0}}$ on local operators (away from the boundary site $12L-$ 1) that appear in the Hamiltonian is then given by
@@
The Pauli operators appearing on the RHS of Eq. (19) are exactly the new local defect terms appearing in Eq. (14). Additionally, the deleted terms from the original no-defect Hamiltonian can also be explained. For example, the transverse field operator $X_{2i_{0}+1}$ is missing in the defect Hamiltonian, which can be explained as follows: after the action of $\widehat{D}_{2i_{0}}$, the original $X_{2i_{0}+1}$ becomes $Z_{2i_{0}+1}Z_{2i_{0}+3}$, while there are no other operators that get mapped to $X_{2i_{0}+1}$. Finally, composing with the domain wall in the coupling space, which exchanges $J$ with $g$ to the right of site $12i_{0}+$ 1, we obtain the duality defect Hamiltonian in Eq. (14).
@@ V.1 Duality defect in the critical TFI model
Figure 2: Diagrammatic presentation of the local terms $\mathcal{O}_{2i}^{(1)}$ and $\mathcal{O}_{2i+1}^{(2)}$ in the context of the fusion tree.
@@
Now that we have recovered the critical TFI Hamiltonian from the fusion category language, we proceed to insert a duality defect, between physical spins at sites $12i_{0}-$ 1 and $12i_{0}+$ 1. States in the duality defect Hilbert space are represented by fusion trees with one vertical branch inserted from below, as illustrated on the LHS in Figure 7.
@@ V.2 Duality defect in the deformed TFI model
Figure 9: Diagrammatic presentation of the local terms $P^{(1)}_{2i+1}$ and $P^{(2)}_{2i}$ in the fusion tree context.
@@
Figure 10: Diagrammatic simplification for the matrix element of $P^{(1)}_{2i+1}$ when $h^{\prime}_{2i+1}\neq h_{2i+1}$.
@@ VI Defect-modified weak integrability breaking
In this Section, we will describe a method to construct defect-modified weak integrability breaking deformations, where the undeformed model is an integrable model with the insertion of an integrable defect. We will demonstrate this method by giving a third way to derive the duality defect Hamiltonian in the KW self-dual deformed TFI model, thanks to the fact that the KW duality defect in the undeformed TFI model is integrable. We also investigate additional defect-modified conserved charges (other than the Hamiltonian) in the TFI model, and confirm their slower decay behavior upon turning on the weak integrability breaking deformation.
@@ VI.1 Weak integrability breaking deformations
The deformed Hamiltonian is a special instance of deformed charges. The $Q_{\alpha}(\lambda)$ and $H(\lambda)$ obtained in this method contain arbitrary long-range interactions whose strength decreases exponentially with the distance. To produce deformations with finite-range interactions, one expands $Q_{\alpha}(\lambda)$ and $H(\lambda)$ as a series in $\lambda$ and performs truncations, say up to order $l-1$, obtaining $Q_{\alpha}^{<l}(\lambda)$ and $H^{<l}(\lambda)$. The cost of doing so, is that $Q_{\alpha}^{<l}(\lambda)$ and $H^{<l}(\lambda)$ don’t strictly commute anymore. Rather, they commute up to a correction at order $\lambda^{l}$. Concretely we have
$\begin{split}\left[Q_{\alpha}^{<l}(\lambda),Q_{\beta}^{<l}(\lambda)\right]&=o(\lambda^{l})~{},\\ \left[Q_{\alpha}^{<l}(\lambda),H^{<l}(\lambda)\right]&=o(\lambda^{l})~{}.\\ \end{split}$ (28)
@@ VI.2 Self-dual deformation in the TFI model
$\begin{split}Q_{\alpha}^{(1)}&=\frac{1}{2}\sum\limits_{j}\{q^{(0)}_{1,j},J^{(0)}_{2\alpha,j}+J^{(0)}_{2\alpha,j+1}\}\\ &-\frac{1}{2}\sum\limits_{j}\{q^{(0)}_{2,j},J^{(0)}_{1\alpha,j}+J^{(0)}_{1\alpha,j+1}\}~{},\end{split}$ (34)
@@
In the following, we will compute the first-order correction to the Hamiltonian $H^{(1)}=-JQ_{2}^{(1)}$ and see that it reproduces the self-dual deformation in the TFI model. We will then proceed to compute $Q_{1}^{(1)}$ and study the decay behavior of the deformed charge $Q_{1}^{<2}(\lambda)=Q_{1}^{(0)}+\lambda Q_{1}^{(1)}$, under the evolution with the KW self-dual deformed Hamiltonian $H^{<2}(\lambda):=H^{(0)}+\lambda H^{(1)}$.
Therefore we have
The first-order deformation to the Hamiltonian, $H^{(1)}$, is given by
$\begin{split}H^{(1)}=&-\frac{J}{2}\sum\limits_{j}\{q^{(0)}_{1,j},J^{(0)}_{22,j}+J^{(0)}_{22,j+1}\}\\ &+\frac{J}{2}\sum\limits_{j}\{q^{(0)}_{2,j},J^{(0)}_{12,j}+J^{(0)}_{12,j+1}\}~{},\end{split}$ (38)
where
Substituing this into Eq. (38), we then have
@@
Up until now, we have been working with an infinite chain. In fact, from Eq. (32), it is clear that the bi-local generator $X^{(0)}=\left[Q_{2}^{(0)}\big{|}Q_{1}^{(0)}\right]$ requires a choice of ordering, which raises potential issues for a finite chain with e.g. periodic boundary condition. On the other hand, the first order deformations $Q_{\alpha}^{(1)}=\text{i}\left[\left[Q_{2}^{(0)}\big{|}Q_{1}^{(0)}\right],Q_{\alpha}^{(0)}\right]$ (as in Eq. (34)) can be expressed in terms of local charge density operators and generalized current density operators, which readily carry over to the case of a periodic chain. In particular, we can apply the resulting deformations such as $H^{(1)}$ in Eq. (40) and $Q_{1}^{(1)}$ in Eq. (41) to our original setup of a periodic chain with $L$ sites. Moreover, we explicitly checked that
@@
We numerically study $A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ for small system sizes through exact diagonalization. After an $o(1)$ initial decay time, $A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ oscillates around a plateau. This is demonstrated in Figure 15. Moreover, similar to [11], we can perform a perturbative analysis for the decay behavior at early times. The details can be found in Appendix B. In particular, the $o(\lambda^{2})$ dependence of $\frac{d}{dt}A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ is expected to vanish, such that its leading-order dependence is at least $o(\lambda^{3})$. In Figure 16, we plot $A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ against $\lambda^{3}t$ for different system sizes and different values of $\lambda$ and observe good data points collapse at early times.
Figure 15: $A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ for $J=1$, $g=0.3$, $L=8,10,12$ and $\lambda=0.1,0.2$.
Figure 16: Short-time decay behavior of $A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ for $J=1$, $g=0.3$, $L=8,10,12$ and $\lambda=0.1,0.2$ respectively. The data points collapse at early times indicates an $o(\lambda^{3})$ leading-order dependence.
@@ VI.3 Weak integrability breaking in presence of KW duality defect
The KW duality defect in the TFI model is an instance of an integrable defect. In the following, we will showcase defect-modified weak integrability breaking using this example. In particular, this gives us a third way to derive the defect Hamiltonian (given by $-JQ_{2}^{D,<2}(\lambda)$) in the deformed TFI model. Additionally, we study the decay of deformed charges in presence of the defect.
@@
As a sanity check, we would like to see that
$\left[Q_{2}^{D,(0)},Q_{1}^{D,(0)}\right]=0~{}.$ (52)
Direct computation yields
Summing these three terms together gives us
Similarly the generalized current density $J_{22}^{D,(0)}$ is
Next we compute the following anti-commutators
together with
@@
Additionally we have

### v4

@@ I Introduction
Compared with open boundaries, long-lived modes associated with defects are much less studied. Recently a special class of defects, the so-called topological defects, have attracted attention from the fields of condensed matter physics, high energy theory, and quantum information [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]. In particular, Ref. [41] considered a 1d transverse field Ising (TFI) model on a circular chain with the insertion of a Kramers-Wannier [67] duality interface and a domain wall in the coupling space, bi-partitioning the chain into a ferromagnetic phase and a paramagnetic phase. In each $\mathbb{Z}_{2}$ charge $\pm 1$ sector of the system, there exists a zero mode supported on neighboring sites around the domain wall, which exactly commutes with the Hamiltonian. More recently, Ref. [55] studied the corresponding problem in a periodically-driven Floquet setting, and found an isolated zero mode localized around the domain wall, exactly commuting with the Floquet unitary. In the special limit where the duality interface is right next to the domain wall thereby shrinking one of the gapped phases to zero size, the zero mode becomes localized around the duality defect. Stability of the zero mode against deformations was probed in [68] for a particular deformed Floquet TFI model. In the presence of the deformation, the zero mode starts to decay. However, the decay mechanism is similar to a reservoir effect, such that for small system sizes the zero mode only partially decays. This can be quantified by the infinite-temperature auto-correlation function of a local operator overlapping with the zero mode, which stabilizes around a finite plateau at infinite time.
@@
Even though our methods apply to more general quantum systems, here we will streamline the ideas using the example of a 1d deformed TFI model in presence of a Kramers-Wannier (KW) duality defect, both in the Hamiltonian setting and in the periodically-driven Floquet setting. First, we derive the deformed defect Hamiltonian given in Eq. (6) for the general case using three different methods: half-chain KW transformation, utilization of techniques from the Ising fusion category, and defect-modified weak integrability breaking deformation. The third method also gives a new framework to study weak integrability breaking in generic integrable systems with integrable defects. The results agree between these three quite distinct methods. We then work out the deformation of defect-modified higher charges in the model, and study their slower decay behavior under the time evolution by the deformed Hamiltonian. Finally, we move on to the corresponding deformed Floquet TFI model with a duality defect. The zero mode localized around the defect starts to decay in the presence of the deformation. Similar to [68], for small system sizes the zero mode only decays partially, as corroborated by a finite plateau of the infinite-temperature auto-correlation function for an overlapping local operator. The main novelty of our work is a framework to study weak integrability breaking in integrable systems with integrable defects. Additionally, we generalized the fusion category construction to include next-to-nearest-neighbor projectors suitable for modeling the deformation.
@@ II Transverse field Ising model with Kramers-Wannier self-dual deformation
The 1d TFI model can be deformed in different ways that are self-dual under the Kramers-Wannier transformation [69, 70]. Here we consider the model first studied in [69]. Our setup is a chain of $L$ sites with site-labels j=1,,Lj=1,...,L. The Hamiltonian for the deformed TFI model with periodic boundary condition is given by
@@ III Duality defect in the self-dual deformed TFI model
Due to the KW self-dual nature of the deformation $H_{\lambda}$, we can construct a duality defect in this deformed TFI model. In later Sections, we will describe three different ways to do so. In Section IV, we construct the duality defect via half-space Kramers-Wannier transformation. In Section V, we illustrate the duality defect construction using techniques from fusion categories, by generalizing previous constructions to next-to-nearest-neighbor projectors that model the deformed theory. In Section VI, we provide another novel way of deriving the duality defect Hamiltonian, as a special kind of weak integrability breaking deformation where the conserved charges in the undeformed model are modified by the defect. The third method gives a new framework to study weak integrability breaking in generic integrable systems with integrable defects, and provides a novel lattice realization of $T\overline{T}$ deformation in the presence of defects. The fact that the results agree between these three different methods serves as a non-trivial check for each individual construction.
@@ III.1 Fusion of two duality defects
In particular, the site $i_{0}-2$ only enters the transformed Hamiltonian through the last three terms in Eq. (12). The Pauli operator $Z_{i_{0}-2}$ commutes with the transformed Hamiltonian, consequently the whole defect Hilbert space admits a decomposition into the $\pm 1$ eigenspace of $Z_{i_{0}-2}$. In the +1+1 eigenspace, Eq. (12) reduces to the no-defect deformed TFI model, albeit on L1L-1 sites labeled by $1,\dots,i_{0}-3,i_{0}-1,\dots,L$. In the 1-1 eigenspace, we obtain the $\eta$-defect Hamiltonian (see Eq. (10)) in the deformed TFI model on L1L-1 sites. This defect Hilbert space decomposition, schematically expressed as
@@ IV Duality defect from half-chain Kramers-Wannier transformation
[…]
@@
The action of $\widehat{D}_{2i_{0}}$ on local operators (away from the boundary site 2L12L-1) that appear in the Hamiltonian is then given by
@@
The Pauli operators appearing on the RHS of Eq. (19) are exactly the new local defect terms appearing in Eq. (14). Additionally, the deleted terms from the original no-defect Hamiltonian can also be explained. For example, the transverse field operator $X_{2i_{0}+1}$ is missing in the defect Hamiltonian, which can be explained as follows: after the action of $\widehat{D}_{2i_{0}}$, the original $X_{2i_{0}+1}$ becomes $Z_{2i_{0}+1}Z_{2i_{0}+3}$, while there are no other operators that get mapped to $X_{2i_{0}+1}$. Finally, composing with the domain wall in the coupling space, which exchanges $J$ with $g$ to the right of site $2i_{0}+1$, we obtain the duality defect Hamiltonian in Eq. (14).
@@ V.1 Duality defect in the critical TFI model
Figure 2: Diagrammatic presentation of the local terms $\mathcal{O}_{2i}^{(1)}$ and $\mathcal{O}_{2i+1}^{(2)}$ in Eq. (22) in the context of the fusion tree.
@@
Now that we have recovered the critical TFI Hamiltonian from the fusion category language, we proceed to insert a duality defect, between physical spins at sites $2i_{0}-1$ and $2i_{0}+1$. States in the duality defect Hilbert space are represented by fusion trees with one vertical branch inserted from below, as illustrated on the LHS in Figure 7.
@@ V.2 Duality defect in the deformed TFI model
Figure 9: Diagrammatic presentation of the local terms $P^{(1)}_{2i+1}$ and $P^{(2)}_{2i}$ in Eq. (24) in the fusion tree context.
@@
Figure 10: Diagrammatic simplification for the matrix element of $P^{(1)}_{2i+1}$ in Eq. (24) when $h^{\prime}_{2i+1}\neq h_{2i+1}$.
@@ VI Defect-modified weak integrability breaking
In this Section, we will describe a method to construct defect-modified weak integrability breaking deformations, where the undeformed model is an integrable model with the insertion of an integrable defect. We will demonstrate this method by giving a third way to derive the duality defect Hamiltonian in the KW self-dual deformed TFI model, thanks to the fact that the KW duality defect in the undeformed TFI model is integrable. We also investigate additional defect-modified conserved charges (other than the Hamiltonian) in the TFI model, and confirm their slower decay behavior upon turning on the weak integrability breaking deformation. We will employ exact diagonalization for small system sizes to study the long time dynamics of infinite-temperature observables, where traditional tensor networks techniques are not well-suited. We refer to Ref. [92] and Ref. [93] for some recent developments of numerical methods to access such regimes of long time dynamics. Interestingly, as we discuss next, the physics of weak integrability breaking deformations is to cause a decay with a higher power of the integrability breaking term. This physics is apparent at short times itself, and therefore exact diagonalization is sufficient to demonstrate it.
@@ VI.1 Weak integrability breaking deformations
The deformed Hamiltonian is a special instance of deformed charges. The $Q_{\alpha}(\lambda)$ and $H(\lambda)$ obtained in this method contain arbitrary long-range interactions whose strength decreases exponentially with the distance. To produce deformations with finite-range interactions, one expands $Q_{\alpha}(\lambda)$ and $H(\lambda)$ as a series in $\lambda$ and performs truncations, say up to order l1l-1, obtaining $Q_{\alpha}^{<l}(\lambda)$ and $H^{<l}(\lambda)$. The cost of doing so, is that $Q_{\alpha}^{<l}(\lambda)$ and $H^{<l}(\lambda)$ don’t strictly commute anymore. Rather, they commute up to a correction at order $\lambda^{l}$. Concretely we have
[…]
@@ VI.2 Self-dual deformation in the TFI model
[…]
@@
We can now apply Eq. (34) to compute first-order deformations. The computational details can be found in Appendix A. Concretely the first-order deformation to the Hamiltonian, $H^{(1)}=-JQ_{2}^{(1)}$, is given by
@@
Rescaling the deformation parameter $\lambda$ by a factor of $4g$, we see that $H^{<2}(\lambda)=H^{(0)}+\lambda H^{(1)}$ indeed reproduces the Hamiltonian for the KW self-dual deformed TFI model. We also remark that, to obtain this result, we have relabeled the summation indices for certain terms, which is justified for the infinite chain setup (as well as for a periodic chain).
We now proceed to compute the deformed charge $Q_{1}^{<2}(\lambda)=Q_{1}^{(0)}+\lambda Q_{1}^{(1)}$, where the undeformed charge $Q_{1}^{(0)}$ is given in Eq. (31). We obtain
@@
Up until now, we have been working with an infinite chain. In fact, from Eq. (32), it is clear that the bi-local generator $X^{(0)}=\left[Q_{2}^{(0)}\big{|}Q_{1}^{(0)}\right]$ requires a choice of ordering, which raises potential issues for a finite chain with e.g. periodic boundary condition. On the other hand, the first order deformations $Q_{\alpha}^{(1)}=\text{i}\left[\left[Q_{2}^{(0)}\big{|}Q_{1}^{(0)}\right],Q_{\alpha}^{(0)}\right]$ (as in Eq. (34)) can be expressed in terms of local charge density operators and generalized current density operators, which readily carry over to the case of a periodic chain. In particular, we can apply the resulting deformations such as $H^{(1)}$ and $Q_{1}^{(1)}$ to our original setup of a periodic chain with $L$ sites. Moreover, we explicitly checked that
@@
We numerically study $A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ for small system sizes through Exact Diagonalization. After an $o(1)$ initial decay time, $A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ oscillates around a plateau. This is demonstrated in the left panel of Figure 15. Moreover, similar to [11], we can perform a perturbative analysis for the decay behavior at early times. The details can be found in Appendix B. In particular, the $o(\lambda^{2})$ dependence of $\frac{d}{dt}A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ is expected to vanish, such that its leading-order dependence is at least $o(\lambda^{3})$. In the right panel of Figure 15, we plot $A_{\infty}^{Q_{1}^{<2}(\lambda)}(t)$ against $\lambda^{3}t$ for different system sizes and different values of $\lambda$ and observe good data points collapse at early times.
Figure 15: Left Panel : The infinite-temperature auto-correlation function $A_{\infty}^{Q_{1}^{<2}(\lambda)}$ as defined in Eq. (39), plotted against time $t$ for $J=1$, $g=0.3$, $L=8,10,12$ and $\lambda=0.1,0.2$. Right Panel : Short-time decay behavior of $A_{\infty}^{Q_{1}^{<2}(\lambda)}$ plotted against $\lambda^{3}t$ for the same set of parameters. The data points collapse at early times indicates an $o(\lambda^{3})$ leading-order dependence.
@@ VI.3 Weak integrability breaking in presence of KW duality defect
The KW duality defect in the TFI model is an instance of an integrable defect. In the following, we will showcase defect-modified weak integrability breaking using this example. In particular, this gives us a third way to derive the defect Hamiltonian (given by $JQ_{2}^{D,<2}(\lambda)$) in the deformed TFI model. Additionally, we study the decay of deformed charges in presence of the defect.
@@
As a sanity check, we can show that
$\left[Q_{2}^{D,(0)},Q_{1}^{D,(0)}\right]=0~,$ (47)
where the details can be found in Appendix A.
Now we are ready to study the deformation generated by the defect-modified bilocal operator $\left[Q_{2}^{D,(0)}\big{|}Q_{1}^{D,(0)}\right]$. Concretely, the first order deformation to the defect Hamiltonian in integrable defect TFI model is
@@
Rescaling $\lambda\to 4g\lambda$ as before, this reproduces the deformation terms, including the extra local modifications, in the defect Hamiltonian Eq. (6) for the self-dual deformed duality-twisted TFI model.
@@
The corresponding $o(\lambda)$ deformation $Q_{1}^{D,(1)}$ is given by (see Appendix A)
@@
where we have again applied the rescaling convention of $\lambda\to 4g\lambda$. Notice that, $Q_{1}^{D,(1)}$ can also be constructed from performing half-space KW transformation on $Q_{1}^{(1)}$ in the no-defect case. We have checked that the results do agree from these two methods.